OpenStudy (anonymous):

If sine of x equals 1 over 2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

jimthompson5910 (jim_thompson5910):

Hint: Pythagorean Trig Identity \[\Large \sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

Ok, so how would I continue off that? May you guide me through the problem please?

jimthompson5910 (jim_thompson5910):

since sine is 1/2, squaring that gives you 1/4 you can replace all of sin^2 with 1/4 like this \[\Large \sin^2(x) + \cos^2(x) = 1\] \[\Large \frac{1}{4} + \cos^2(x) = 1\]

jimthompson5910 (jim_thompson5910):

now isolate cos(x)

OpenStudy (anonymous):

we subract 1/4 from both sides

OpenStudy (anonymous):

so now we square root both sides right?

jimthompson5910 (jim_thompson5910):

correct on both statements

OpenStudy (anonymous):

then how do we get tangent

jimthompson5910 (jim_thompson5910):

\[\Large \tan(x) = \frac{\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

Ok, hold on. Let me put everything together.

OpenStudy (anonymous):

cos(x) = \[\sqrt{.75}\] is that how it should be written

jimthompson5910 (jim_thompson5910):

or \[\Large \sqrt{\frac{3}{4}} = \frac{\sqrt{3}}{\sqrt{4}} = \frac{\sqrt{3}}{2}\]

jimthompson5910 (jim_thompson5910):

most books I've seen use \[\Large \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

oh ok, so \[\tan(x) = \frac{ 1 }{ 2 } /\frac{ \sqrt{3} }{ 2 }\]

jimthompson5910 (jim_thompson5910):

Yes correct. Now simplify

OpenStudy (anonymous):

How would I do that?

jimthompson5910 (jim_thompson5910):